OpenStudy (anonymous):

solve: 2x^2 = -7x + 15

OpenStudy (kkutie7):

\[2x^{2}=-7x+15\] move everything to one side \[2x^{2}+7x-15=0\] you can try and factor this, but because I'm lazy I'll just plug it into the quadratic formula \[a=2,b=7,c=-15\]

OpenStudy (kkutie7):

\[\frac{ -b \pm \sqrt{b^{2}-4(a)(c)} }{ 2(a) }\]

OpenStudy (kkutie7):

\[x=-5,\frac{3}{2}\]